# CHIMERA: Adaptive Cache Injection and Semantic Anchor Prompting for Zero-shot Image Morphing with Morphing-oriented Metrics

### v2

@@ 4 Proposed Method: CHIMERA
Overall Pipeline. As shown in Fig. 4, Given input images AA and $B$, we first project them into the latent space using a DDIM inversion (DDIM) [43] to obtain $z_{A}=\texttt{DDIM}(A)$ and $z_{B}=\texttt{DDIM}(B)$. We then perform spherical interpolation (slerp) to form the $K$-morphing latent images:
@@
During DDIM inversion [52] to obtain the inverted latents for AA and $B$, we cache multi-scale U-Net features. Specifically, we record the down, mid, and up features as:
@@ 4.1 Adaptive Cache Injection (ACI)
Previous image morphing methods [24, 27, 7] typically interpolate only the latents obtained from DDIM and then perform denoising. However, prior state–of–the–art (SOTA) method [7] provides limited input–aligned guidance to the diffusion model, often producing images that deviate from both AA and $B$. To address this limitation, we propose Adaptive Cache Injection (ACI), which guides the denoising process by injecting depth- and timestep-adaptive cached multi–scale features.
As described in Sec. 4, during DDIM, we cache down ($\mathbf{D}$), mid ($\mathbf{M}$), and up ($\mathbf{U}$) features for each input (see Eq. 3). Then, for each morphing index $k$, cached features from AA and $B$ are blended via Eq. 4.
@@
where $F_{S}$ denotes the denoising diffusion feature at layer $S$, and $\tilde{F}_{S}$ represents the feature after adding the cached feature as a residual. Through this overall process, ACI provides layer–wise, aligned timestep–aware guidance, enabling the generation of smooth and faithful morphing results aligned with both AA and $B$.
@@ 4.3 Global-Local Consistency Score (GLCS)
Proposed Metric. We propose Global-Local Consistency Score to evaluate morphing quality with two complementary factors. First, the Global Consistency Score (GCS) measures domain consistency. It checks whether each image follows the expected global trend between the two input images AA and $B$. We obtain this trend by interpolating the endpoint similarities with slerp, so the sequence should change in a balanced way from AA to $B$. Second, the Local Consistency Score (LCS) measures smoothness. It checks whether the similarity of each image changes smoothly with respect to its neighbors. Thus, LCS captures local continuity along the morphing transition. We use a DiffSim-based [45] bounded similarity $s(\cdot,\cdot)$, which is sensitive to low-level structure and also reflects style and semantic similarity. Both GCS and LCS are clamped to $[0,1]$ for stability and interpretability. GLCS combines these two perspectives and is high only when the sequence is globally well-mixed and locally smooth:
@@ 5.2 Qualitative Evaluations
To demonstrate the effectiveness of our proposed CHIMERA, we provide a qualitative comparison with existing methods in Fig. 6. As shown by the red arrows in Fig. 6 (a), the transition is relatively natural near the input images AA and $B$, but it becomes increasingly abrupt toward the center of the sequence. As observed in the red arrows of Fig. 6 (b), DiffMorpher loses semantic information from both AA and $B$ when the input pair becomes more dissimilar, leading to degraded perceptual quality. In Fig. 6 (c), FreeMorph achieves good perceptual quality and smoothness but still introduces saturated colors and over-smoothing, producing textures that deviate from the input pair. In contrast, CHIMERA preserves perceptual quality while maintaining the semantics of both AA and $B$, achieving smooth transitions. Notably, in Fig. 6 (d), the red arrows show that our model retains the stone texture on the person’s arm and coherently blends the person’s face into the rocky mountain. Additional visual comparisons are provided in the Suppl.
@@ Appendix A Extended Experiment Results
This section provides additional quantitative and qualitative evaluations that supplement Sec. 5.1 and Sec. 5.2. In Sec. A.1, we present qualitative and quantitative results for the setting where 14 morphing images are generated between input images AA and $B$. Unlike Sec. 5.1, which reports quantitative results for the 5-image morphing setting, this section evaluates CHIMERA under a longer morphing transition to assess the general applicability of the proposed method. In addition, Sec. A.2 provides further qualitative results for the 5-frame morphing scenario discussed in Sec. 5.2.
@@ A.2 Additional Qualitative Result on 5-Image Morphing
Fig. 8 and Fig. 9 present qualitative results for the setting where five morphing images are generated between AA and $B$. As shown in panels (a) and (e) of Fig. 8 and Fig. 9 (red arrows), IMPUS produces frames with abrupt transitions. In panels (b) and (f), the morphing images exhibit good smoothness, but the red arrows highlight collapsed structures or noticeable artifacts. In panels (c) and (g), the transitions remain smooth, yet the red arrows indicate a tendency toward excessively saturated colors. In contrast, panels (d) and (h) of Fig. 8 and Fig. 9 show that the proposed CHIMERA preserves both domain consistency and smoothness.
@@ Appendix C Ablation Study on Inversion-Denoising Timestep Mapping (IDM)
When the inversion timesteps are fixed, Table 6 shows that our IDM-based model (a) achieves the best quantitative performance. As illustrated in Fig. 13, fixing inversion to early, mid, or late produces undesired results: the model generates images that deviate from the input images AA and $B$ (Fig. 13 (b), (c)), or produces structurally unstable results with severe artifacts (Fig. 13 (d)). In each case, the red arrows in the figure explicitly indicate the regions where these degradations occur.
@@ Appendix E Detail Decription of GLCS
Let AA and $B$ be the endpoint images, and let $\{I_{k}\}_{k=1}^{K}$ be the predicted morphing images ordered from AA to $B$. We adopt a DiffSim-based bounded similarity [45], denoted by
@@
where $\alpha_{k}$ encodes the ideal mixing ratio between the two endpoints AA and $B$.
@@ E.1 Effects of GCS and LCS
Figure 15: Qualitative examples showing how the GCS component of GLCS aligns with human perception. Blue arrows indicate frames where the domains of AA and $B$ are properly mixed according to the perceived interpolation ratio, while red arrows indicate frames where the two domain cues are not well reflected given the same interpolation ratio.
@@
Fig. 15 reports the effect of GCS on selected morphing images. In Fig. 15 (a) and (b), the red lines and dots indicate cases with low GCS scores, while the blue lines and dots indicate cases with high GCS scores. In Fig. 15 (a), the Morphing-0 image is highly similar to image AA and also shares a similar background with Morphing-1, resulting in a high GCS score of 90.789. In contrast, Morphing-1 should strongly reflect the wolf and moderately reflect the human from image AA, but it fails to do so, leading to a low GCS score. Moreover, Morphing-2 does not properly reflect either the wolf or the human, and thus shows the lowest score among the morphing images (Fig. 15 (b) shows a similar case). Unlike (a) and (b), panels (c) and (d) exhibit consistently high GCS values across the morphing sequence, and human observers also perceive strong domain consistency that includes both domains of AA and $B$. This indicates that (c) and (d) have higher domain consistency than (a) and (b). These results demonstrate that the proposed GCS can evaluate domain consistency in a manner that aligns well with human perception.
@@ E.2 Comparison between traditional metric and GLCS
Fig. 17 provides a qualitative comparison between $\mathrm{FID}_{\text{local}}$ and GCS. As shown in Fig. 17, the first rows of (a) and (b) achieve better $\mathrm{FID}_{\text{local}}$ scores than the second rows. However, visual inspection reveals that the third image in the first row of (a) does not properly include both domains of AA and $B$, and the fourth image even produces a result that is unrelated to image $B$. Similarly, in the first row of (b), the third and fourth images contain almost no elements from image $B$. These observations indicate that $\mathrm{FID}_{\text{local}}$ does not align well with human perception when evaluating domain consistency, since it only compares the overall distributions of $A,B$ and the morphing images.
In contrast, the proposed GCS evaluates whether each image properly reflects both domains of AA and $B$ according to the interpolation ratio. As a result, the second rows of (a) and (b), which better preserve domain consistency, are assigned higher quality scores than the first rows. This demonstrates that GCS provides a more human-aligned assessment of domain consistency.

### v4

@@ 4 Proposed Method: CHIMERA
Overall Pipeline. As shown in Fig. 4, Given input images $A$ and $B$, we first project them into the latent space using a DDIM inversion (DDIM) [41] to obtain $z_{A}=\texttt{DDIM}(A)$ and $z_{B}=\texttt{DDIM}(B)$. We then perform spherical interpolation (slerp) to form the $K$-morphing latent images:
@@
During DDIM inversion [50] to obtain the inverted latents for $A$ and $B$, we cache multi-scale U-Net features. Specifically, we record the down, mid, and up features as:
@@ 4.1 Adaptive Cache Injection (ACI)
Previous image morphing methods [52, 55, 7] typically interpolate only the latents obtained from DDIM and then perform denoising. However, prior state–of–the–art (SOTA) method [7] provides limited input–aligned guidance to the diffusion model, often producing images that deviate from both $A$ and $B$. To address this limitation, we propose Adaptive Cache Injection (ACI), which guides the denoising process by injecting depth- and timestep-adaptive cached multi–scale features.
As described in Sec. 4, during DDIM, we cache down ($\mathbf{D}$), mid ($\mathbf{M}$), and up ($\mathbf{U}$) features for each input (see Eq. 3). Then, for each morphing index $k$, cached features from $A$ and $B$ are blended via Eq. 4.
@@
where $F_{S}$ denotes the denoising diffusion feature at layer $S$, and $\tilde{F}_{S}$ represents the feature after adding the cached feature as a residual. Through this overall process, ACI provides layer–wise, aligned timestep–aware guidance, enabling the generation of smooth and faithful morphing results aligned with both $A$ and $B$.
@@ 4.3 Global-Local Consistency Score (GLCS)
Proposed Metric. We propose Global-Local Consistency Score to evaluate morphing quality with two complementary factors. First, the Global Consistency Score (GCS) measures domain consistency. It checks whether each image follows the expected global trend between the two input images $A$ and $B$. We obtain this trend by interpolating the endpoint similarities with slerp, so the sequence should change in a balanced way from $A$ to $B$. Second, the Local Consistency Score (LCS) measures smoothness. It checks whether the similarity of each image changes smoothly with respect to its neighbors. Thus, LCS captures local continuity along the morphing transition. We use a DiffSim-based [43] bounded similarity $s(\cdot,\cdot)$, which is sensitive to low-level structure and also reflects style and semantic similarity. Both GCS and LCS are clamped to $[0,1]$ for stability and interpretability. GLCS combines these two perspectives and is high only when the sequence is globally well-mixed and locally smooth:
@@ 5.2 Qualitative Evaluations
To demonstrate the effectiveness of our proposed CHIMERA, we provide a qualitative comparison with existing methods in Fig. 6. As shown by the red arrows in Fig. 6 (a), the transition is relatively natural near the input images $A$ and $B$, but it becomes increasingly abrupt toward the center of the sequence. As observed in the red arrows of Fig. 6 (b), DiffMorpher loses semantic information from both $A$ and $B$ when the input pair becomes more dissimilar, leading to degraded perceptual quality. In Fig. 6 (c), FreeMorph achieves good perceptual quality and smoothness but still introduces saturated colors and over-smoothing, producing textures that deviate from the input pair. In contrast, CHIMERA preserves perceptual quality while maintaining the semantics of both $A$ and $B$, achieving smooth transitions. Notably, in Fig. 6 (d), the red arrows show that our model retains the stone texture on the person’s arm and coherently blends the person’s face into the rocky mountain. Additional visual comparisons are provided in the Suppl.
@@ Appendix A Extended Experiment Results
This section provides additional quantitative and qualitative evaluations that supplement Sec. 5.1 and Sec. 5.2. In Sec. A.1, we present qualitative and quantitative results for the setting where 14 morphing images are generated between input images $A$ and $B$. Unlike Sec. 5.1, which reports quantitative results for the 5-image morphing setting, this section evaluates CHIMERA under a longer morphing transition to assess the general applicability of the proposed method. In addition, Sec. A.2 provides further qualitative results for the 5-frame morphing scenario discussed in Sec. 5.2.
@@ A.2 Additional Qualitative Result on 5-Image Morphing
Fig. 8 and Fig. 9 present qualitative results for the setting where five morphing images are generated between $A$ and $B$. As shown in panels (a) and (e) of Fig. 8 and Fig. 9 (red arrows), IMPUS produces frames with abrupt transitions. In panels (b) and (f), the morphing images exhibit good smoothness, but the red arrows highlight collapsed structures or noticeable artifacts. In panels (c) and (g), the transitions remain smooth, yet the red arrows indicate a tendency toward excessively saturated colors. In contrast, panels (d) and (h) of Fig. 8 and Fig. 9 show that the proposed CHIMERA preserves both domain consistency and smoothness.
@@ Appendix C Ablation Study on Inversion-Denoising Timestep Mapping (IDM)
When the inversion timesteps are fixed, Table 6 shows that our IDM-based model (a) achieves the best quantitative performance. As illustrated in Fig. 13, fixing inversion to early, mid, or late produces undesired results: the model generates images that deviate from the input images $A$ and $B$ (Fig. 13 (b), (c)), or produces structurally unstable results with severe artifacts (Fig. 13 (d)). In each case, the red arrows in the figure explicitly indicate the regions where these degradations occur.
@@ Appendix E Detail Decription of GLCS
Let $A$ and $B$ be the endpoint images, and let $\{I_{k}\}_{k=1}^{K}$ be the predicted morphing images ordered from $A$ to $B$. We adopt a DiffSim-based bounded similarity [43], denoted by
@@
where $\alpha_{k}$ encodes the ideal mixing ratio between the two endpoints $A$ and $B$.
@@ E.1 Effects of GCS and LCS
Figure 15: Qualitative examples showing how the GCS component of GLCS aligns with human perception. Blue arrows indicate frames where the domains of $A$ and $B$ are properly mixed according to the perceived interpolation ratio, while red arrows indicate frames where the two domain cues are not well reflected given the same interpolation ratio.
@@
Fig. 15 reports the effect of GCS on selected morphing images. In Fig. 15 (a) and (b), the red lines and dots indicate cases with low GCS scores, while the blue lines and dots indicate cases with high GCS scores. In Fig. 15 (a), the Morphing-0 image is highly similar to image $A$ and also shares a similar background with Morphing-1, resulting in a high GCS score of 90.789. In contrast, Morphing-1 should strongly reflect the wolf and moderately reflect the human from image $A$, but it fails to do so, leading to a low GCS score. Moreover, Morphing-2 does not properly reflect either the wolf or the human, and thus shows the lowest score among the morphing images (Fig. 15 (b) shows a similar case). Unlike (a) and (b), panels (c) and (d) exhibit consistently high GCS values across the morphing sequence, and human observers also perceive strong domain consistency that includes both domains of $A$ and $B$. This indicates that (c) and (d) have higher domain consistency than (a) and (b). These results demonstrate that the proposed GCS can evaluate domain consistency in a manner that aligns well with human perception.
@@ E.2 Comparison between traditional metric and GLCS
Fig. 17 provides a qualitative comparison between $\mathrm{FID}_{\text{local}}$ and GCS. As shown in Fig. 17, the first rows of (a) and (b) achieve better $\mathrm{FID}_{\text{local}}$ scores than the second rows. However, visual inspection reveals that the third image in the first row of (a) does not properly include both domains of $A$ and $B$, and the fourth image even produces a result that is unrelated to image $B$. Similarly, in the first row of (b), the third and fourth images contain almost no elements from image $B$. These observations indicate that $\mathrm{FID}_{\text{local}}$ does not align well with human perception when evaluating domain consistency, since it only compares the overall distributions of $A,B$ and the morphing images.
In contrast, the proposed GCS evaluates whether each image properly reflects both domains of $A$ and $B$ according to the interpolation ratio. As a result, the second rows of (a) and (b), which better preserve domain consistency, are assigned higher quality scores than the first rows. This demonstrates that GCS provides a more human-aligned assessment of domain consistency.